Accessibility Assistance

Skip to Content

Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
March 29, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Howell, Michael Jay, Tom Minton, Patricia Ruwell, Josh Spurgin, Evan Wack

Guest: Rachel Schipper
  1. Bobbie Meng is retiring as of 30 April.
  2. Cynthia and her dog won the dog/owner dress-alike contest in High Springs. They placed second in the dog/owner look-alike contest.
  3. College basketball March Madness banter.
  4. Staff Directory - please verify that your information is accurate. Will to forward the email announcement to IT staff.
  5. Staff evaluations are done. Faculty evaluations need to be done. Updated position descriptions need to be done this week.
  6. MSL IT offices - there are power issues. We are working with Facilities to get it fixed. It may not be possible.
Review Minutes of 3/22/10
  1. Nets # 2 - remove two sentences.
Ongoing IT Updates

  1. PO for the 174 workstations has gone out. Evan would like to get the estimated date of delivery.
  2. Altiris - the new Statement of Work was sent to Purchasing.
  1. Tapes in vault - movement, but need indication by name as to which tapes can be disposed of or need to be saved. John Ingram is reviewing the request from Joe to retain 474 tapes and the costly option of transferring them to a more readable format.
  2. Eastside - waiting for new building diagram. Logan got an email from Mary Byrd regarding wiring. We don't know the number of ports needed. The offices are configured with a standard layout.
  1. Public printing - now that Bobbie is retiring, it will be handed off to Evan.
  2. ReturnIL/Clean Slate - Clean Slate is not working on install. Josh is testing it. We may have to put more time into ReturnIL. The management of Clean Slate is better, but it is not working now.
  3. Help Desk - document devised to assist students entering tickets. Trish and Evan to schedule a meeting with Telecom. Working on ideas for the all staff email to announce the implementation of the new IT phones and Help Desk.
  4. Anti-virus is working great. HIP and Site advisor have helped to curb infections.
  5. Videoconferencing - Thursday they are coming to install new microphones. There is a communication problem they will also fix.
  6. Win7 - Staff is done. Public - log on registry edit does not work. We are working on it.
  7. Altiris - waiting for PO.
  8. Time-Limited workstations - approved. Need to get new quote, then proceed with order.
  9. Request for two new phones in MSL. Need to try to re-provision existing phones.
  1. Patron Update failed again. They say the version of driver we use to access that data is out of date. It would be labor intensive to update. We decided to put it on the new server that Logan created for us.
  2. ARL Stats - ongoing.
  3. Library Hours GUI - ready for review by Tom. Need to check with Misty's group.
Network/Server Administration
  1. VT calendar - just about done. It is the first app on the DHCP server.
  2. DHCP - starting to progress. We settled that Altiris is doable. Need to request exactly what we need to CNS. Need to decide if we will give back some of our IP ranges. Should be ready by the beginning of summer.
  3. CNS virtualization status - SQL server done. Linus app server done. Web space available. DLC space available. Hope to have DLC done by today. GatorLink mailbox storage shift in a week or 10 days.
  4. CNS backup - we are backing up every weekend. DLC Production SAN - the last week's data is available, but not the current ongoing week. Laurie is okay with it.
  5. SharePoint pushed to June. Allows for time to upgrade to SharePoint 2010, the most current version. Standard template? Need to decide. Probably a good idea. Staff will need training (as will we). Need to get with Misty regarding Town Hall Meeting dates, as Judy is traveling and she will want to be there.
  1. ALEPH training ongoing.
  2. Doug's training is helpful.
Web Projects
  1. Proxy - Cliff is working on it. Has a test version with Shibboleth compatibility. Need a test server. The look of the log-in page will change.
  2. CMS - problematic figuring out the plug-ins and widgets.
  3. Open Access project - needs to be given to the Dean by Thursday - electronic signatures for the Institutional Repository. Funding from UF. There is a legal service that acts as a signature.
  4. Regional Depository site for GovDocs - sortable database. State plan and timeline.
  5. Survey Monkey has expired. Sonia to renew.
DLC General Issues
  1. JPEG 2K server - next to go virtual. Will be converting DLC all this week.